Consider the Following Reaction : 2Mg + O2 -> 2MgO
Determine how many grams of MgO will be formed if 40.0g of each reactant is used?
1
Expert's answer
2017-12-12T11:57:07-0500
m(Mg)=40.0g
m(O2)=40.0g
m(MgO) - ?
2Mg + O2 -> 2MgO
n(Mg)=m(Mg)/M(Mg)=40g / 24g/mol =1.67mol

n(O2)=m(O2)/M(O2)=40g / 32g/mol = 1.25mol

2n(Mg)<n(O2) => n(MgO) = n(Mg)=1.67mol
m(MgO)= n(MgO)*M (MgO)=1.67mol * 40g/mol = 66.8g